# Energy-Entropy Regularization: The True Power of Minimal Looped Transformers

### v1

@@ 3.1 Tsallis Entropy
Tsallis entropy provides a natural extension of this classical entropy measure. Given $w$ possible outcomes with probabilities $p_{1},\ldots,p_{w}$, the Tsallis entropy is defined as
@@ 5.3 Kinetic Regularization
To enforce asymptotic stability within the iterative loop, we introduce a kinetic penalty. Unlike standard $L_{2}$ weight regularization, we penalize the squared Euclidean norm of the state residuals. This term acts as a dissipative force (damping) that encourages the system to reach an equilibrium:
[…]
By varying the rollout depth $T\in[16,64]$, we encourage the discovery of length-invariant attractors rather than transient paths. This process undergoes a structural phase transition: early liquid exploration of the manifold is gradually replaced by crystallization as the penalty induces stability in the latent trajectory.
@@ B.4 Remark
$Z_{k+1}=Z_{k}+\alpha(H_{q})\!\bigl(S(X+Z_{k})\bigr)\bigl(F(X+Z_{k})-Z_{k}\bigr).$
where $\alpha$ is the update step size. The theorem suggest a straight forward entropy control $H_{q}$. One can find a suitable $H_{q}$ easily following the following properties:
  • Positivity: $0<\alpha(H_{q})\leq 1$.
  • Monotonicity: $H_{q}^{(1)}\leq H_{q}^{(2)}\implies\alpha(H_{q}^{(1)})\leq\alpha(H_{q}^{(2)})$. Low entropy implies smaller step.
  • Contraction: There must exists $\beta(H_{q})$ such that $\alpha(H_{q})\beta(H_{q})<1$.
  • Satuation at High Entropy: Since we do not want to slow down the convergence unneccesarily, so when the attention is spreaded, $\alpha(H_{q})\approx 1$.

### v2

@@ 3.1 Tsallis Entropy
Tsallis entropy provides a natural extension of classical entropy measure. Given $w$ possible outcomes with probabilities $p_{1},\ldots,p_{w}$, the Tsallis entropy is defined as
@@ 5.3 Kinetic Regularization
To enforce asymptotic stability within the iterative loop, we introduce a kinetic penalty. Unlike standard $L_{2}$ weight regularization, we penalize the Euclidean displacement of the state residuals. This term acts as a dissipative force (damping) that encourages the system to reach an equilibrium:
$\mathcal{L}_{\text{Kinetic}}=\frac{1}{2BL}\sum_{b=1}^{B}\sum_{\ell=1}^{L}|Z_{T,b\ell}-Z_{0,b\ell}|_{2}$ (3)
where $B$ is the batch size and $L$ is the sequence length. By varying the input length $L\in[16,64]$ during training, we encourage the discovery of length-invariant attractors rather than transient paths. This process undergoes a structural phase transition: early liquid exploration of the manifold is gradually replaced by crystallization as the penalty induces stability in the latent trajectory.
@@ B.4 Remark
$Z_{k+1}=Z_{k}+\alpha(S_{q})\bigl(F(X+Z_{k})-Z_{k}\bigr).$
where $\alpha(\cdot)$ is the update step size. The theorem suggest a straight forward entropy control $\alpha(S_{q})$. One can find a suitable $\alpha(S_{q})$ easily following the following properties:
  • Positivity: $0<\alpha(S_{q})\leq 1$.
  • Monotonicity: $S_{q}^{(1)}\leq S_{q}^{(2)}\implies\alpha(S_{q}^{(1)})\leq\alpha(S_{q}^{(2)})$. Low entropy implies smaller step.
  • Contraction: There must exists $\beta(S_{q})$ such that $\alpha(S_{q})\beta(S_{q})<1$.
  • Satuation at High Entropy: Since we do not want to slow down the convergence unneccesarily, so when the attention is spreaded, $\alpha(S_{q})\approx 1$.